ROTATIONAL SPECTROSCOPIC INVESTIGATIONS OF CH_{4} - H_{2}S COMPLEX

\underline{AISWARYA LAKSHMI P.} and E. ARUNAN, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore 560012, India.

Recently, Raghavendra and Arunan [1] have shown that methane can form a hydrogen bonded complex with HX (X = F, Cl, OH, SH), in which H − X points towards the carbon of methane. Microwave spectroscopy has confirmed the structure of CH_{4} − HF/HCl/H_{2}O complexes. No experimental data is available for CH_{4} - H_{2}S complex to the best of our knowledge. Pulsed Nozzle Fourier Transform Microwave Spectrometer has been used to study the rotational spectrum of CH_{4} - H_{2}S complex. Two progressions were observed. Both could be fitted independently to a linear top. The ground state rotational constant is determined to be B = 2683.100(1)MHz and the distortion constant to be D_{J} = 0.09413(9)MHz. Based on this rotational constant, the intermolecular separation is determined to be 4.136Å. The other state has a rotational constant B = 2593.05(1)MHz and a negative distortion constant, D_{J} = −0.0089(7)MHz. The negative distortion constant implies rotational - vibrational coupling and the progression arises from some excited internal rotor/torsional state. Experiments are in progress to look for more states and isotopologues of this complex. Details will be presented in the talk.

1.Raghavendra et al.Chem. Phys. Lett. 467(2008)37.